What differences would you expect to see in the physical properties of [Mn(CO)6]2+ and [MnBr6]-4 and why?
Answer #1

Both [Mn(CO)6]2+ and [MnBr6]4- has Mn2+ state which is 4s0 3d5.

But CO is strong field ligand and Br is weak field liegand.

hence when d orbitals is splitted into t2g and eg .

we see pairing of electrons in t2g when CO is present . Hence we get 1 lone pair of electron.

Magnetic moment = sqrt ( n( n+2) = sqrt ( 1( 3) = 1.732

for [Mn[Br6]4- complex we see that all 5 d orbitals are half filled.

Hence magnetic moment = sqrt ( 5 ( 5+2)) = 5.9

Thus magnteic moment of [MnBr6]4- is larger than [Mn(CO)6]2+

thus in magntic field we can spot the different

Another visible property is colour.

There are spin allowed transitions for [Mn[CO]6}2+ i.e d5 in low spin state.

But there are no spin allowed d-d transitiosn for [Mn[Br6]4- ,

hence Mn[CO]6]2+ is darker in colour compoared to [Mn[Br]6]4- which si pale or colourless.